OpenStudy (cwrw238):

There are 5 keys. Each one will open one of five cars. If i pick one of the keys at random and try to open one of the cars what is the probability that its the right key?

OpenStudy (anonymous):

1/5 ?

OpenStudy (cwrw238):

right - you guys are pretty good - most people say its 1/25

OpenStudy (cwrw238):

- it doesn't matter which key you pick

OpenStudy (unklerhaukus):

\[\langle x\rangle=\frac{1/5}{1}=\frac 15\]